# IndicSentEval: How Effectively do Multilingual Transformer Models encode Linguistic Properties for Indic Languages?

### v1

@@ Abstract
Transformer-based models have revolutionized the field of natural language processing. To understand why they perform so well and to assess their reliability, several studies have focused on questions such as: Which linguistic properties are encoded by these models, and to what extent? How robust are these models in encoding linguistic properties when faced with perturbations in the input text? However, these studies have mainly focused on BERT and the English language. In this paper, we investigate similar questions regarding encoding capability and robustness for 8 linguistic properties across 13 different perturbations in 6 Indic languages, using 9 multilingual Transformer models (7 universal and 2 Indic-specific). To conduct this study, we introduce a novel multilingual benchmark dataset, IndicSentEval, containing approximately $\sim$47K sentences. Surprisingly, our probing analysis of surface, syntactic, and semantic properties reveals that while almost all multilingual models demonstrate consistent encoding performance for English, they show mixed results for Indic languages. As expected, Indic-specific multilingual models capture linguistic properties in Indic languages better than universal models. Intriguingly, universal models broadly exhibit better robustness compared to Indic-specific models, particularly under perturbations such as dropping both nouns and verbs, dropping only verbs, or keeping only nouns.Overall, this study provides valuable insights into probing and perturbation-specific strengths and weaknesses of popular multilingual Transformer-based models for different Indic languages. We make our code and dataset publicly available [^1] https://tinyurl.com/IndicSentEval.
@@ 1 Introduction
Transformer-based language models (Vaswani et al., 2017), pretrained in both mono-lingual and multilingual contexts using millions of text documents, have demonstrated substantial enhancements in the performance of various natural language processing (NLP) tasks (Kenton and Toutanova, 2019; Pires et al., 2019; Conneau et al., 2020; Radford et al., 2019; Brown et al., 2020; Wang et al., 2019, 2018). To understand what types of linguistic properties (surface, syntactic and semantic) are encoded across layers of Transformer-based models, initial studies have investigated the layer-wise representations via a range of probing tasks Adi et al. (2017); Hupkes et al. (2018); Conneau et al. (2018); Rogers et al. (2020); Jawahar et al. (2019); Mohebbi et al. (2021). However, these studies focus on solely on English. Although, several studies have examined the presence of shared representation aspects across widely spoken languages in multilingual models (Chi et al., 2020; Acs et al., 2023), unfortunately, there is no work that investigates to what extent multilingual Transformer-based models encode linguistic properties for different languages. Further, recent neuro-AI studies have revealed that the brain uses different parsing strategies for different linguistic properties, which further differ across languages Zhang et al. (2022); Oota et al. (2023). This inspired us to study the nuances of how multilingual Transformer-based models capture linguistic properties across layers and languages.
@@
Indic languages offer a rich tapestry of linguistic features that contribute significantly to global linguistic diversity. Therefore, in this paper, we focus on Indic languages. We study the extent to which the linguistic properties of Indic languages are captured by two kinds of multilingual models: universal and Indic-specific models. Universal multilingual models have been pretrained using data which contains a small and varying fraction of Indic languages across languages and models (Table 13 in Appendix). In contrast, Indic-specific models (Kakwani et al., 2020; Dabre et al., 2021; Kumar et al., 2022) have been trained exclusively on Indic language data only. Specifically, we focus on 6 Indic languages: three Indo-European languages (Hindi, Marathi, Urdu) and three Dravidian languages (Telugu, Kannada, Malayalam).
@@
Even if multilingual models effectively encode linguistic properties of Indic languages, such encodings may not be robust to input text perturbations for particular models. A lack of robustness may render some models less reliable than others for real-world applications. Although there exist many such studies Wang et al. (2021); Jin et al. (2020); Li et al. (2020); Garg and Ramakrishnan (2020); Sanyal et al. (2022); Neerudu et al. (2023) on robustness analysis of Transformer-based models, they focus on English, and downstream tasks. Robustness analysis for the prediction of linguistic properties in a multilingual setting for Indic languages is under explored. Hence, we systematically study how various multilingual Transformer-based models may differ in their robustness for different linguistic properties with respect to different kinds of input text perturbations across languages.
@@
We analyze 7 universal multilingual language models, each pretrained on data spanning $\sim$100+ languages, but containing only a small amount of Indic language data. These models include mBERT Pires et al. (2019), XLM-R Conneau et al. (2020), InfoXLM Chi et al. (2021), mGPT Shliazhko et al. (2024), XGLM Lin et al. (2022), BLOOM Scao et al. (2022), and mT5 Xue et al. (2021). Additionally, we examine 2 Indic-specific models, IndicBERT Kakwani et al. (2020) and MuRIL Khanuja et al. (2021), which are trained using corpora for Indic languages along with English. While it is expected that these Indic-specific models would be better at encoding and exhibit greater robustness for Indic language input, we pose several questions: Are there some linguistic properties that are better encoded by universal models? Are the universal models more robust to particular types of perturbations? How effectively and robustly are English language properties encoded by these universal and Indic-specific models?
@@
To perform such detailed analyses, we curate a novel multilingual dataset, IndicSentEval, from resources generated by the “Indian Language Machine Translation” (ILMT) initiative [^2] https://tinyurl.com/6hhdcce9. This dataset contains information about three types of linguistic properties for each Indic language: surface, syntactic and semantic. To investigate encoding and robustness ability of multilingual models, we design probing tasks for prediction of each property. The surface task probes whether the model learns a representation that is predictive of sentence length (SentLen). Syntactic tasks test for sensitivity to word order i.e. bigram shift (BShift) and the depth of the syntactic tree (TreeDepth). Semantic tasks check for the subject and the direct object number in the main clause (SubjNum and ObjNum, respectively). The tasks mentioned so far have also been discussed in English-focused studies. However, morphology of Indic languages is significantly different from that of English, primarily regarding the main verb used in a sentence. This prompts us to expand our analysis to encompass three additional semantic probing tasks related to the main verb within the sentence: verb gender (VerbGen), verb number (VerbNum), and verb person (VerbPer). Fig. 1 shows the conceptual diagram of our study.
Our work leads to the following surprising results. For the English language, our probing analysis of surface, syntactic, and semantic properties reveals that almost all multilingual models demonstrate consistent encoding performance and a similar language hierarchy. However, we observe mixed results for Indic languages.Specifically, Indic-specific multilingual models capture the language hierarchy better for Indic languages, while universal models show inconsistent results. Intriguingly, universal models broadly exhibit better robustness compared to Indic-specific models, with the most impactful perturbations involving dropping both nouns and verbs, dropping only verbs, or keeping only nouns.
Overall, the main contributions of this paper are as follows.
  • We conduct an extensive study on how 9 multilingual Transformer-based models capture 8 linguistic properties across 6 Indic languages.
  • We contribute a novel dataset, IndicSentEval, with $\sim$47K sentences across the 6 languages.
  • We find that Indic-specific models like MuRIL and IndicBERT best capture linguistic properties for Indic languages, while universal models like mBERT, InfoXLM, mGPT and BLOOM show mixed results across properties.
@@
  • Surprisingly, our robustness analysis with respect to 13 text perturbations shows that universal multilingual models (InfoXLM, BLOOM, mGPT, XGLM and mT5) demonstrate higher resilience to perturbations compared to BERT-like models (mBERT, IndicBERT and MuRIL).
2 Related Work
Our work is most closely related to that of Jawahar et al. (2019) and Mohebbi et al. (2021), who focus on understanding the interpretability of language models via probing and observing linguistic structures in English. Our study on multilingual language models for six Indic languages complements their English-focused research.
Our work also relates to the growing literature that creates resources for low-resource Indic languages. Several recent studies have developed resources including Indic language models Khan et al. (2024), the Indic NLP suite Kakwani et al. (2020), Indic BART Dabre et al. (2021), Indic NLG Kumar et al. (2022), IndicMT Eval Dixit et al. (2023), and Naamapadam Mhaske et al. (2023). Our research is the first to interpret both universal and Indic-specific models for Indic languages through probing representations and assessment of robustness, while also providing implications for future research directions. Overall, we complement these works by studying the linguistic structures of a wide range of low-resource Indic languages using multilingual models.
3 Probing Tasks
Probing tasks (Adi et al., 2017; Hupkes et al., 2018; Jawahar et al., 2019; Mohebbi et al., 2021) help unpack the linguistic features that may be encoded in neural language models. These probing tasks are formulated as prediction tasks and focus on various aspects of sentence structure. We experiment with eight probing tasks to evaluate how effectively multilingual models encode linguistic properties across six Indic languages: Hindi (hi), Telugu (te), Marathi (mr), Kannada (kn), Urdu (ur), and Malayalam (ml). These eight probing tasks are grouped into three categories: surface (SentLen), syntactic (BShift and TreeDepth), and semantic (SubjNum, ObjNum, VerbGender, VerbPerson and VerbNumber). We selected these tasks because they cover different aspects of language and require varying levels of abstraction and generalization. These tasks involve 3 binary and 5 multi-class classification problems. The specifics of the initial five probing tasks are thoroughly outlined in Conneau et al. (2018) as well. We also provide brief descriptions for each probing task in the following, with a summary of class labels for each task in Table 1.
Surface level tasks (1) Sentence Length (SentLen): Here, the objective is to predict the number of words in sentences, which has been grouped into 8 categories as shown in Table 1.
@@ 4 IndicSentEval Dataset
We curate the IndicSentEval dataset from resources generated by the ILMT initiative, which serves as an Indic language counterpart to SentEval Conneau et al. (2018) and offers labeled data for the eight probing tasks. We utilize the morph and chunk level Indic languages data Tandon and Sharma (2017); Bhatt et al. (2009); Xia et al. (2008) available in Shakti Standard Format (SSF) Bharati et al. (2007, 1995). SSF is a highly readable representation for storing Indic language data with linguistic annotations. Fig. 4 in Appendix shows an example of a Hindi sentence in SSF format.
Task ||||C|||| Labels
@@
Table 1: Probing task details: number of classes and class labels.
IndicSentEval curation details. For each property, we gather data per language as follows. (1) SentLen: We iterate through all the nodes in the SSF format representation of the sentence and count number of words in each chunk. (2) TreeDepth: We utilize the data from the dependency tree to perform a traversal, specifically employing breadth first search. This traversal enables us to calculate the tree depth of the sentence. (3) BShift: For this task, we randomly (probability=0.2) select the sentences from the dataset, and then a randomly selected bigram (equal probability for all bigrams) is inverted. Sentences with inverted bigrams are marked as 1, and the rest as 0. (4) SubjNum/ObjNum: We identify the subject/object (a noun that can be singular or plural) of a sentence using the assigned semantic roles in the SSF format, and use the NN/NNS annotations. (5) Verb Gender/Person/Number: We first locate the chunk containing the main verb using the annotated chunk label from the SSF format. Then, we extract the gender/person/number information from the annotated morph output. Examples for each probing tasks per language are shown in detail in Table 11 in the Appendix. Also, detailed statistics of number of samples across 6 Indic languages for 8 probing tasks are provided in Appendix Tables 9 and 10.
@@ 5 Text Perturbation Analysis
Probing tasks can help us understand the extent to which various multilingual models encode different linguistic properties. But which particular words (such as nouns, verbs and others) in the input sentence help models encode these structures? To answer this question, we experiment with three different categories of perturbations: AppendR, DropText and Positional. We chose these perturbations because they simulate types of noise found in real datasets by introducing different degrees of noise and variation in the input text. Particularly, we experiment with the following text perturbations.
@@
DropText. DropText perturbations reflect situations where critical information is missing or only certain types of words are retained, which is common in incomplete or corrupted datasets. This includes, DropNV (dropping words based on their part-of-speech tag, specifically both nouns (N) and verbs (V)), DropN (dropping all nouns), DropV (dropping all verbs), DropRN (dropping one random noun), DropRV (dropping one random verb), KeepNV (dropping all words except nouns and verbs), KeepN (dropping all words except nouns), and KeepV (dropping all words except verbs). DropText perturbations are designed to provide deeper insights into word-level attention mechanisms within models, specifically aiming to determine whether models focus more on objects (nouns), actions (verbs), or the contextual elements surrounding these key parts-of-speech.
@@
Overall, these text perturbations help in understanding the contribution of specific word types and sentence structures to the encoding capabilities of multilingual models. Tables 14-19 in the Appendix display examples of text perturbations for each language.
Figure 2: Probing task results: Layerwise accuracy comparisons between various multilingual representations on surface (top row) and syntactic (bottom two rows) probing tasks. We report the layerwise probing accuracies for individual multilingual models in Appendix Figs. 5 to 10.
@@ 6 Methodology
Multilingual language models. We experiment with nine multilingual Transformer-based models (details in Table 12 in Appendix). First seven have been trained across 100+ languages; IndicBERT and MuRIL support 12 and 17 Indic languages, respectively. Representations are extracted from the encoder layers of mBERT-base, IndicBERT-base, mT5-base, XLM-R, InfoXLM and MuRIL; and from the decoder layers of BLOOM, mGPT and XGLM. We use pretrained model checkpoints from HuggingFace (Wolf et al., 2020).
Probing tasks classifier. To evaluate each probing task using a multilingual model representation, we use logistic regression Wright (1995) classifier with sentence representations as input and the probing task label as target. The base model is frozen. We use mean pooling across tokens to get the sentence representation.
@@
Hyper-parameters. We train logistic regression with a regularization parameter $C$=20 and use a L2 penalty term. For multi-class tasks, we use the “multinomial” setting while training. All experiments were done on a machine with a T4 GPU.
Evaluation metrics. Similar to earlier studies Conneau et al. (2018); Jawahar et al. (2019); Mohebbi et al. (2021), for all the probing tasks, we use classification accuracy as evaluation metric. Let $A_{c}$ and $A_{p}$ be accuracy of a model on the clean and perturbed test sets, respectively. To evaluate the perturbation results for probing tasks, we use robustness score defined as $1-\frac{A_{c}-A_{p}}{A_{c}}$. Typically, robustness score of a model ranges between 0 and 1 where 0 indicates that the model is not robust, and 1 indicates that the model is completely robust. Score > 1 suggests that the model’s performance improves when the perturbation is applied.

### v2

@@ Abstract
Transformer-based models have revolutionized the field of natural language processing. To understand why they perform so well and to assess their reliability, several studies have focused on questions such as: Which linguistic properties are encoded by these models, and to what extent? How robust are these models in encoding linguistic properties when faced with perturbations in the input text? However, these studies have mainly focused on BERT and the English language. In this paper, we investigate similar questions regarding encoding capability and robustness for 8 linguistic properties across 13 different perturbations in 6 Indic languages, using 9 multilingual Transformer models (7 universal and 2 Indic-specific). To conduct this study, we introduce a novel multilingual benchmark dataset, IndicSentEval, containing approximately $\sim$47K sentences. Our probing analysis of surface, syntactic, and semantic properties reveals that, while almost all multilingual models demonstrate consistent encoding performance for English, surprisingly, they show mixed results for Indic languages. As expected, Indic-specific multilingual models capture linguistic properties in Indic languages better than universal models. Intriguingly, universal models broadly exhibit better robustness compared to Indic-specific models, particularly under perturbations such as dropping both nouns and verbs, dropping only verbs, or keeping only nouns. Overall, this study provides valuable insights into probing and perturbation-specific strengths and weaknesses of popular multilingual Transformer-based models for different Indic languages. We make our code and dataset publicly available [^1] https://github.com/aforakhilesh/IndicBertology.
@@ 1 Introduction
Transformer-based language models (Vaswani et al., 2017), pretrained for both mono-lingual and multilingual contexts using millions of text documents, have demonstrated substantial enhancements in the performance of various natural language processing (NLP) tasks (Kenton and Toutanova, 2019; Pires et al., 2019; Conneau et al., 2020; Radford et al., 2019; Brown et al., 2020; Wang et al., 2019, 2018). To understand what types of linguistic properties (surface, syntactic and semantic) are encoded across layers of Transformer-based models, initial studies have investigated the layer-wise representations via a range of probing tasks (Adi et al., 2017; Hupkes et al., 2018; Conneau et al., 2018; Rogers et al., 2020; Jawahar et al., 2019; Mohebbi et al., 2021). However, these studies focus solely on English. Although, several studies have examined the presence of shared representation aspects across widely spoken languages in multilingual models (Chi et al., 2020; Acs et al., 2024), unfortunately, there is no work that investigates the extent to which multilingual Transformer-based models encode linguistic properties for different Indic languages. Further, recent neuro-AI studies have revealed that the brain uses different parsing strategies for different linguistic properties, which further differ across languages (Zhang et al., 2022; Oota et al., 2023). This inspired us to study the nuances of how multilingual Transformer-based models capture linguistic properties across layers and languages.
@@
Indic languages offer a rich tapestry of linguistic features that contribute to the global linguistic diversity. Hence, in this paper, we focus on Indic languages. We study the degree to which linguistic properties of Indic languages are captured by two kinds of multilingual models: universal and Indic-specific models. Universal multilingual models have been pretrained using a variety of pretraining objectives and using data which contains a small and varying fraction of Indic languages across languages and models (Tables 1920, and 21 in Appendix F). Indic-specific models (Kakwani et al., 2020; Dabre et al., 2021; Kumar et al., 2022) have been specifically trained on Indic language data only. Specifically, we focus on 6 Indic languages: three Indo-European languages (Hindi, Marathi, Urdu) and three Dravidian languages (Telugu, Kannada, Malayalam).
@@
Even if multilingual models encode linguistic properties of Indic languages effectively, such encodings may not be robust to input text perturbations for particular models. Lack of such robustness may make some models less reliable than others for real-world applications. Although there exist many such studies (Wang et al., 2021; Jin et al., 2020; Li et al., 2020; Garg and Ramakrishnan, 2020; Sanyal et al., 2022; Neerudu et al., 2023) on robustness analysis of Transformer-based models, they focus on English, and on downstream tasks. Robustness analysis for prediction of linguistic properties in a multilingual setting for Indic languages is under explored. Hence, we systematically study how various multilingual Transformer-based models may differ in their robustness for different linguistic properties with respect to different kinds of input text perturbations across languages. We provide detailed related work in Appendix A, focusing on three aspects: (i) probing in non-English and multilingual contexts, (ii) differences from English-centric findings, and (iii) critiques of probing methodology and recent advances.
@@
We analyze 7 universal multilingual language models, each pretrained on data spanning $\sim$100+ languages, with only a small amount of Indic language data. These include mBERT (Pires et al., 2019), XLM-R (Conneau et al., 2020), InfoXLM (Chi et al., 2021), mGPT (Shliazhko et al., 2024), XGLM (Lin et al., 2022), BLOOM (Scao et al., 2022), and mT5 (Xue et al., 2021). Additionally, we examine 2 Indic-specific models, IndicBERT (Kakwani et al., 2020) and MuRIL (Khanuja et al., 2021), which are trained using corpora for Indic languages along with English. While it is expected that these Indic-specific models would be better at encoding and robustness for Indic language input, are there some linguistic properties which are better encoded by universal models? Are the universal models more robust to particular perturbation types? How effectively and robustly are English language properties encoded by these universal and Indic-specific models?
@@
To perform such detailed analyses, we curate a novel multilingual dataset, IndicSentEval, from resources generated by the “Indian Language Machine Translation” (ILMT) initiative. This dataset contains information about three types of linguistic properties per Indic language: surface, syntactic and semantic. To investigate encoding and robustness ability of multilingual models, we design probing tasks for prediction of each property. The surface task probes whether the model learns a representation which is predictive of sentence length (SentLen). Syntactic tasks test for sensitivity to word order, i.e., bigram shift (BShift) and the depth of the syntactic tree (TreeDepth). Semantic tasks check for the subject and the direct object number in the main clause (SubjNum and ObjNum, respectively). Tasks mentioned so far were discussed in English focused studies too. However, morphology for Indic languages is significantly different from English, primarily in regard to the main verb used in a sentence. This prompts us to expand our analysis to encompass three additional semantic probing tasks related to the main verb within the sentence: verb gender (VerbGen), verb number (VerbNum), and verb person (VerbPer). Fig. 1 shows the conceptual diagram of our study.
@@
Overall, the main contributions of this paper are as follows. (1) We perform an extensive study of the degree to which 9 multilingual Transformer-based models capture 8 linguistic properties across 6 Indic languages. (2) We contribute a novel dataset, IndicSentEval, with $\sim$47K sentences across the 6 languages. (3) We find that Indic-specific models like MuRIL and IndicBERT best capture linguistic properties for Indic languages, while universal models like mBERT, InfoXLM, mGPT and BLOOM show mixed results across properties. (4) Surprisingly, our robustness analysis with respect to 13 text perturbations shows that universal multilingual models (InfoXLM, BLOOM, mGPT, XGLM and mT5) demonstrate higher resilience to perturbations compared to BERT-like models (mBERT, IndicBERT and MuRIL).
@@ 2 IndicSentEval Dataset
We curate the IndicSentEval dataset from resources generated by the ILMT initiative, which serves as an Indic language counterpart to SentEval (Conneau et al., 2018) and offers labeled data for the eight probing tasks. We utilize the morph and chunk level Indic languages data (Tandon and Sharma, 2017; Bhatt et al., 2009; Xia et al., 2008) available in Shakti Standard Format (SSF) (Bharati et al., 2007, 1995). SSF is a highly readable representation for storing Indic language data with linguistic annotations. Fig. 4 in Appendix B shows an example of a Hindi sentence in SSF format.
Probing Tasks. Probing tasks (Adi et al., 2017; Hupkes et al., 2018; Jawahar et al., 2019; Mohebbi et al., 2021) help unpack the linguistic features possibly encoded in neural language models. These probing tasks are formulated as prediction tasks and focus on several aspects of sentence structure. We experiment with eight probing tasks to evaluate how effectively multilingual models encode linguistic properties across six Indic languages: Hindi (hi), Telugu (te), Marathi (mr), Kannada (kn), Urdu (ur), and Malayalam (ml). These eight probing tasks are grouped into three categories: surface (SentLen), syntactic (BShift and TreeDepth), and semantic (SubjNum, ObjNum, VerbGen, VerbPerson and VerbNumber). We selected these tasks because they cover different aspects of language and require different levels of abstraction and generalization. These tasks involve 3 binary and 5 multi-class classification problems. The specifics of the initial five probing tasks are thoroughly outlined in (Conneau et al., 2018) as well. We also provide brief descriptions for each probing task in Appendix D, with a summary of class labels for each task in Table 1.
Task ||C|| Labels
@@
Table 1: Probing task details: number of classes ($|C|$) and class labels.
IndicSentEval curation details. For each property, we gather data per language as follows. (1) SentLen: We iterate through all the nodes in the SSF format representation of the sentence and count number of words in each chunk. (2) TreeDepth: We utilize the data from the dependency tree to perform a traversal, specifically employing breadth first search. This traversal enables us to calculate the tree depth of the sentence. (3) BShift: For this task, we randomly (probability=0.2) select the sentences from the dataset, and then a randomly selected bigram (equal probability for all bigrams) is inverted. Sentences with inverted bigrams are marked as 1, and the rest as 0. (4) SubjNum/ObjNum: We identify the subject/object (a noun that can be singular or plural) of a sentence using the assigned semantic roles in the SSF format, and use the NN/NNS annotations. (5) Verb Gender/Person/Number: We first locate the chunk containing the main verb using the annotated chunk label from the SSF format. Then, we extract the gender/person/number information from the annotated morph output. Detailed statistics of number of samples across 6 Indic languages for 8 probing tasks are provided in Tables 10 and 11 in Appendix C. Further, Table 12 in Appendix D shows examples for each probing task per language.
@@ 3 Text Perturbation Analysis
While probing reveals what linguistic features are present in representations, perturbation-based robustness tests address a different question: “how robust is a model to various types of noise and whether it can still understand and process the core meaning of the text despite the introduced variation?” Since our current IndicSentEval dataset lacks noise examples, hence, in this study, we conducted our perturbation analysis on the input dataset to evaluate its robustness. To answer this question, we experiment with three different categories of perturbations: AppendR, DropText and Positional. We chose these perturbations because they simulate types of noise found in real datasets by introducing different degrees of noise and variation in the input text. Particularly, we experiment with the following text perturbations.
@@
DropText. DropText perturbations reflect situations where critical information is missing or only certain types of words are retained, which is common in incomplete or corrupted datasets. This includes DropNV (dropping words based on their part-of-speech tag, specifically both nouns (N) and verbs (V)), DropN (dropping all nouns), DropV (dropping all verbs), DropRN (dropping one random noun), DropRV (dropping one random verb), KeepNV (dropping all words except nouns and verbs), KeepN (dropping all words except nouns), and KeepV (dropping all words except verbs). DropText perturbations are designed to provide deeper insights into word-level attention mechanisms within models, specifically aiming to determine whether models focus more on objects (nouns), actions (verbs), or the contextual elements surrounding these key parts-of-speech.
@@
Overall, these text perturbations help in understanding the contribution of specific word types and sentence structures to the encoding capabilities of multilingual models. Tables 13-18 in Appendix E display examples of perturbations for each language.
Figure 2: Probing task results: Layerwise accuracy comparisons between various multilingual representations on surface (top row) and syntactic (bottom two rows) probing tasks. We report the layerwise probing accuracies for individual multilingual models in Figs. 5 to 10 in Appendix G.
@@ 4 Methodology
Multilingual language models. We experiment with nine multilingual Transformer-based models (listed in Table 19 in Appendix F). First seven have been trained across 100+ languages; IndicBERT and MuRIL support 12 and 17 Indic languages, respectively. Representations are extracted from the encoder layers of mBERT-base, IndicBERT-base, mT5-base, XLM-R, InfoXLM and MuRIL; and from the decoder layers of BLOOM, mGPT and XGLM. We use pretrained model checkpoints from Hugging Face (Wolf et al., 2020).
Probing tasks classifier. To evaluate each probing task using a multilingual model representation, we use logistic regression (Wright, 1995) classifier with sentence representations as input and the probing task label as target. The base model is frozen. We use mean pooling across tokens to get the sentence representation. Details of the hyperparameters are reported in Appendix F.
@@
Evaluation metrics. Similar to earlier studies (Conneau et al., 2018; Jawahar et al., 2019; Mohebbi et al., 2021), for all the probing tasks, we use classification accuracy as the evaluation metric. Let $A_{c}$ and $A_{p}$ be accuracy of a model on the clean and perturbed test sets, respectively. To evaluate the perturbation results for probing tasks, we use robustness score (RS) defined as $RS=1-\frac{A_{c}-A_{p}}{A_{c}}$. Typically, RS of a model ranges between 0 and 1 where 0 indicates that the model is not robust, and 1 indicates that the model is completely robust. Score > 1 suggests that the model’s performance improves when the perturbation is applied.